OpenStudy (anonymous):

lim as x approches infinity of 7-x+8x^2/4x^3+2

OpenStudy (xishem):

\[\lim_{x \rightarrow \infty}\frac{8x^2-x+7}{4x^3+2}\]Since this is of indeterminate form ∞/∞, you can use L Hopital's rule: \[\lim_{x \rightarrow \infty}\frac{16x-1}{12x^2}=\lim_{x \rightarrow \infty}\frac{16}{24x}=0\]

OpenStudy (anonymous):

Thanks!